# Adaptive Visuo-Tactile Fusion with Predictive Force Attention for Dexterous Manipulation

### v1

@@ I INTRODUCTION
However, vision and touch exhibit fundamentally different characteristics. Vision is more global and provides contextual information, while tactile sensing is more local and offers precise feedback on physical contact. Effectively integrating these distinct sensory streams into a coherent understanding remains a significant challenge. Therefore, we ask the question: how can we design multimodal robotic policy frameworks that effectively address and bridge this cross-modal heterogeneity between vision and touch, leveraging their complementary strengths?
@@
In addition, researchers have also explored feature-level fusion [9, 10, 2, 11]. These works primarily encourage the network to fully leverage both unique and shared information from each modality to form a more comprehensive representation of the current state. However, they overlook that different modalities require varying levels of attention at different manipulation stages [12], and in some cases, additional modalities can even be a distraction [1, 13]. Predicting contact probability to modulate sensory features is one strategy to address such distractions. FoAR [13], for instance, implements this concept by using its contact predictions to adjust force features from end-effector force/torque in real-time, helping to avoid noise interference. However, it relies on a predefined threshold to label contact and assumes that vision features always dominate. In contrast, the force signal evolves with physical interaction, capturing contact state and interaction progression, and provides a natural guidance signal for attention modulation across modalities without task-specific assumptions or manual labels.
In this work, we propose Adaptive Visuo-Tactile Fusion with Predictive Force Attention (AdapTac), a method designed to adaptively integrate visual and tactile modalities for dexterous manipulation. Specifically, 1) AdapTac introduces a force-guided attention fusion module, where force signals serve as queries, and visual and tactile features act as keys and values. By computing attention weights between the force and different modalities, our policy adaptively adjusts the attention of visual and tactile, avoiding task-specific labeling requirements and improving generalization to different tasks. 2) AdapTac also incorporates a self-supervised auxiliary loss, which uses a diffusion model-based force head to predict future force signals during training. This auxiliary future force prediction reinforces tactile representations and alleviates data imbalance. The predicted future forces, combined with observed force information, provide temporal context of contact to effectively guide visuo-tactile fusion, ensuring context-aware modality weighting and efficient attention adjustment.
@@ II RELATED WORK
Imitation Learning (IL) enables robots to learn policies by imitating expert demonstrations [14, 15, 16]. Recent advances in policy architectures include the Diffusion Policy [17], which uses diffusion models [18] to capture diverse action modes, and the Action Chunk Transformer (ACT) [19], which predicts structured action chunks for coherent long-term behavior. Meanwhile, input representations have shifted from RGB images [17, 20, 21] toward 3D geometric forms [7, 22, 23, 24]. Early attempts with multi-view voxelized point clouds [25, 20] faced high computational costs and challenging calibration, leading to single-view methods like the 3D Diffusion Policy [22]. However, its PointNet-based encoder [26] showed limitations for dexterous tasks. More point cloud encoders, including Transformer variants [27] and PointNet++ [28], still struggle with noise and sparsity issues. RISE [29] tackles this using sparse convolutional structures [30], achieving better performance on real-world tasks. Despite these developments, most IL methods mainly leverage visual information and overlook important tactile feedback, which is crucial for precise manipulation. Our work incorporates both 3D vision and tactile feedback for dexterous manipulation.
@@ IV METHOD
Problem Statement: We focus on the problem of fusing tactile data from distributed tactile sensors with point cloud data to learn adaptive visuo-tactile dexterous manipulation policies. Given a sequence of observations $\{\mathcal{O}_{t-h+1},\ldots,\mathcal{O}_{t}\}$ over a history horizon $h$, where each observation $\mathcal{O}_{t}=\{\mathcal{O}_{t}^{{pc}},\mathcal{O}_{t}^{{tac}}\}$ consists of point cloud and tactile inputs, our objective is to learn a policy $\pi$ that predicts the next $n$ robot actions $A_{t}=\{{A_{t+1},A_{t+2},\ldots,A_{t+n}}\}$. Each action includes the robot’s translation, rotation, and hand joint positions.
@@ IV-A Force-Guided Attention Fusion
To align feature dimensions, we apply separate MLPs to the point cloud and tactile features: $e^{{pc}}=\mathbf{g}_{{pc}}(\mathcal{Z}^{{pc}})$ and $e^{{tac}}=\mathbf{g}_{tac}(\mathcal{Z}^{{tac}})$. Based on these aligned features, we obtain the query, key, and value representations as follows:
@@
This module encourages the policy to adaptively adjust the weights of visual and tactile features based on the force at different manipulation stages. Moreover, this design allows the policy to prioritize tactile over visual when necessary, rather than considering vision as the dominant modality all the time [13].
@@ IV-B Future Force Prediction and Guidance
To address this issue, we design a self-supervised future force prediction task during training. We employ a transformer-based diffusion force head to predict the future net force $\mathbf{F}^{n}_{p}$. The diffusion force head takes the visual and tactile features as input and predicts the future net force $\mathbf{F}^{n}_{p}$ at the next $n$ steps. The predicted future force $\mathbf{F}^{n}_{p}$ is then concatenated with the observed net force $\mathbf{F}^{n}_{\mathcal{O}}$ to form a guide force $\mathbf{F}^{n}_{g}$. The updated query is computed as:
$Q^{\mathbf{F}}=g_{F}([\mathbf{F}^{n}_{\mathcal{O}},\mathbf{F}^{n}_{p}])W_{Q}$ (4)
where $g_{F}$ is an MLP that projects the force vectors into the query space, and $W_{Q}$ is a learnable projection matrix.
The guide force $\mathbf{F}^{n}_{g}$ is then used as the query for the force-guided attention fusion module, as shown in Fig. 2. This design allows the attention module to leverage both current and future contact information, enabling better adjustments of the weights of visual and tactile modalities.
@@ IV-C Visuo-Tactile Policy Learning
We integrate the force-guided attention fusion, future force prediction, and guidance mechanism into an imitation learning framework for visuo-tactile dexterous manipulation. Specifically, we adopt RISE [29] as our base policy architecture, which is a 3D diffusion model-based policy. The extracted visual and tactile features are first used for future force prediction, and the predicted future force $\mathbf{F}^{n}_{p}$ is concatenated with the observed net force $\mathbf{F}^{n}_{\mathcal{O}}$ to form the guide force $\mathbf{F}^{n}_{g}$. This guide force is then used as the query in the cross-modal attention module. Finally, the fused features are passed to the diffusion action head [17] to predict the future action. During the training, the policy loss $\mathcal{L}{\pi}$ combines with an additional future force prediction loss $\mathcal{L}_{\text{ffp}}$ to update the network. The total loss is defined as:
@@ V-C Manipulation Policy Comparison
As shown in Tab. I, our proposed method (Ours) outperforms all baselines across all tasks. The vision-only baseline, RISE, performs well on tasks with strong visual information (Open Box, Reorientation) but struggles significantly on the Flip task. This task requires precise manipulation driven by tactile feedback, which RISE lacks. This observation aligns with findings reported in 3DTacDex [3]. The 3DTacDex-P baseline, which employs visuo-tactile concatenation, demonstrates poor overall performance, even underperforming RISE. We also observed that during training, 3DTacDex-P achieves lower hand-joint prediction error compared to other methods. These are likely due to overfitting to tactile patterns, which are similar across different expert demonstrations, while visuals varied a lot due to the larger workspace compared to the original method. The FoAR baseline, despite careful threshold selection for contact prediction, is unstable. While it performs well on the Reorientation task, it fails on both Open Box and Flip tasks. In these failures, the policy reaches the object but cannot apply the correct manipulation forces. Analysis of the expert dataset shows frequent contact changes during manipulation. Manual thresholding struggles to handle these variations, resulting in inconsistent contact labels and poor performance. In contrast, Ours mitigates tactile overfitting by dynamically adjusting visuo-tactile attention across diverse tasks without relying on manual labeling, demonstrating robust generalization and consistent performance.
@@ V-E Importance of Future Force Prediction and Guidance
As shown in Tab. III, without both FFP and FFG, the success rate is only 50%. Incorporating observed force prediction (OFP) and guidance (OFG) increases the success rate to 70%, demonstrating the effectiveness of force prediction and guidance. Replacing observed force (OFP) with future force prediction (FFP) further raises the success rate to 90%, similar to Ours, highlighting the importance of predicting future force. We also observe that for both OFP and FFP, more attention is given to the tactile modality upon contact, indicating that force prediction effectively reinforces tactile modality and improves balance. However, Ours w/o FFG often requires multiple attempts before succeeding, leading to an AEL of 166, whereas Ours typically succeeds on the first attempt, achieving an AEL of only 113. Additionally, Ours w/o FFG exhibits more risky behaviors, such as continuously squeezing the board. These findings underscore the importance of FFG for efficient and robust dexterous manipulation.
@@ V-F Analysis of Attention
We visualize the attention weights at different steps across all tasks, as shown in Fig. 3. For all tasks, during the reaching stages, more attention is given to the visual features. As the hand manipulates the object, the attention shifts toward the tactile features. We also observe that, especially in the reorientation task, which involves multiple instances of reaching and reorientation, Ours adjusts attention weights appropriately as the stage changes. This suggests that our policy effectively distributes attention to different modalities at various stages to complete the tasks.
@@ V-G Generalization on Unseen Objects
Ours 75% 63% 75% 71%
To validate the generalization of our method, we conduct experiments on all tasks using various objects with different colors and geometries, as shown in Fig. 4. Each object is tested twice, with a different random pose each time. As shown in Tab. IV, our policy achieves an average success rate of 71% on unseen objects. Even for objects with significantly different geometries from the training set, such as a small paper cup in the reorientation task or a small whiteboard eraser in the flipping task, our method demonstrates strong generalization.

### v2

@@ I INTRODUCTION
However, vision and touch exhibit fundamentally different characteristics. Vision is more global and provides contextual information, while tactile sensing is more local and offers precise feedback on physical contact. Effectively integrating these distinct sensory streams into a coherent understanding remains a significant challenge. Therefore, we ask the question: how can we design multisensory robotic policy frameworks that effectively bridge the heterogeneity between vision and touch, leveraging their complementary strengths?
@@
In addition, researchers have also explored feature-level fusion [9, 10, 2, 11]. These works primarily encourage the network to fully leverage both unique and shared information from each modality to form a more comprehensive representation of the current state. However, they overlook that different modalities require varying levels of attention at different manipulation stages [12], and in some cases, additional modalities can even be a distraction [1, 13]. Predicting contact probability to modulate sensory features is one strategy to address such distractions. FoAR [13], for instance, implements this concept by using its contact predictions to adjust force features from end-effector force/torque in real-time, helping to avoid noise interference. However, it relies on a predefined threshold to label contact and assumes that vision features always dominate, which limits its scalability.
In this work, we propose Adaptive Visuo-Tactile Fusion with Predictive Force Attention (AdapTac), a method for adaptively integrating visual and tactile modalities in dexterous manipulation. Unlike prior approaches that rely on fixed assumptions (e.g., vision dominance) or manually defined thresholds, AdapTac leverages contact-induced force that naturally reflects contact states and interaction dynamics. Specifically, 1) AdapTac introduces a force-guided attention module, where force signals serve as queries, and visual/tactile features as keys and values. By computing attention weights between the force and different modalities, our policy adaptively adjusts the attention of visual and tactile, eliminating the need for task-specific labels and improving generalization. 2) AdapTac also incorporates a self-supervised auxiliary loss, which uses a diffusion force head to predict future force signals during training. This auxiliary future force prediction reinforces the tactile modality and alleviates data imbalance. The predicted future forces, combined with observed force information, provide temporal context of contact to effectively guide visuo-tactile fusion, ensuring context-aware modality weighting and efficient attention adjustment.
@@ II RELATED WORK
Imitation Learning (IL) allows robots to learn policies by imitating expert demonstrations [14, 15, 16]. Recent advancements include the Diffusion Policy [17], which uses diffusion models [18] for diverse actions, and the Action Chunk Transformer (ACT) [19], which predicts structured action chunks for long-term behavior. Additionally, the shift from RGB images [17, 20, 19] to 3D geometric forms [7, 21, 22, 23, 24] captures richer spatial information, improving the model’s generalization. Early works with multi-view voxelized point clouds [25, 26] laid the foundation for spatial understanding. More recently, the 3D Diffusion Policy [21] has further highlighted the potential of point clouds for action policies. However, such encoders like DP3 [21], PointNet++ [27], and Transformer variants models [28] still struggle with noise and sparsity, limiting their real-world performance. RISE [22] improves robustness with sparse convolutional network [29]. Despite these advances, most IL methods rely heavily on visual data, overlooking the critical role of tactile feedback for precise manipulation. Our approach integrates both 3D vision and tactile feedback to enhance dexterous manipulation.
@@ IV METHOD
Problem Statement: We focus on the problem of fusing tactile data from distributed tactile sensors with point cloud data to learn adaptive visuo-tactile dexterous manipulation policies. Given a sequence of observations $\{\mathcal{O}_{t-h+1},\ldots,\mathcal{O}_{t}\}$ over a history horizon $h$, where each observation $\mathcal{O}_{t}=\{\mathcal{O}_{t}^{{pc}},\mathcal{O}_{t}^{{tac}}\}$ consists of point cloud $\mathcal{O}_{t}^{{pc}}\in\mathbb{R}^{N\times 3}$ and tactile inputs $\mathcal{O}_{t}^{{tac}}\in\mathbb{R}^{120\times 12}$ [3], our objective is to learn a policy $\pi$ that predicts the next $n$ robot actions $A_{t}=\{{A_{t+1},A_{t+2},\ldots,A_{t+n}}\}$. Each action $A_{t}\in\mathbb{R}^{25}$ consists of robot’s translation $\mathbf{t}\in\mathbb{R}^{3}$, rotation $\mathbf{r}\in\mathbb{R}^{6}$, and hand joint positions $\mathbf{q}\in\mathbb{R}^{16}$.
@@ IV-A Force-Guided Attention Fusion
To align the feature dimensions, we apply separate MLPs to the point cloud and tactile features: $e^{{pc}}=\mathbf{g}_{{pc}}(\mathcal{Z}^{{pc}})\in\mathbb{R}^{512},\,e^{{tac}}=\mathbf{g}_{tac}(\mathcal{Z}^{{tac}})\in\mathbb{R}^{512}$. Based on these aligned features, we obtain the query, key, and value representations as follows:
@@
This module enables adaptive weighting of visual and tactile features based on force, allowing the policy to prioritize touch when needed rather than always relying on vision [13].
@@ IV-B Future Force Prediction and Guidance
To address this issue, we design a self-supervised future force prediction task during training, where we introduce a transformer-based [52] diffusion head [18] for future net force prediction. The diffusion force head takes the visual and tactile features as input and predicts the future net force $\mathbf{F}^{n}_{p}$ at the next $n$ steps.
We then concatenate the observed net force $\mathbf{F}^{n}_{\mathcal{O}}$ and predicted future forces $\mathbf{F}^{n}_{p}$ to form the guide force $\mathbf{F}^{n}_{g}$, and project it as the query:
$Q^{\mathbf{F}}=g_{F}([\mathbf{F}^{n}_{\mathcal{O}},\mathbf{F}^{n}_{p}])W_{Q}.$ (4)
where $g_{F}$ is an MLP that projects the force vectors into the query space, and $W_{Q}$ is a learnable projection matrix. This query guides the attention module using both current and future contact information.
@@ IV-C Visuo-Tactile Policy Learning
We integrate the force-guided attention fusion and future force prediction into an imitation learning framework for visuo-tactile dexterous manipulation. Specifically, we adopt RISE [22] as our base policy architecture, which is a 3D diffusion model-based policy. The extracted visual and tactile features are first used for future force prediction, and the predicted future force $\mathbf{F}^{n}_{p}$ is concatenated with the observed net force $\mathbf{F}^{n}_{\mathcal{O}}$ to form the guide force $\mathbf{F}^{n}_{g}$. This guide force is then used as the query in the cross-modal attention module. Finally, the fused features are passed to the diffusion action head [17] to predict the future action. During the training, the policy loss $\mathcal{L}{\pi}$ combines with an additional future force prediction loss $\mathcal{L}_{\text{ffp}}$ to update the network. The total loss is defined as:
@@ V-C Manipulation Policy Comparison
As shown in Tab. I, our proposed method (Ours) outperforms all baselines across all tasks. The vision-only baseline, RISE, performs well on tasks with strong visual information (Open Box, Reorientation) but struggles significantly on the Flip task. This task requires precise manipulation driven by tactile feedback, which RISE lacks. This observation aligns with findings reported in 3DTacDex [3]. The 3DTacDex-P baseline, which employs visuo-tactile concatenation, demonstrates poor overall performance, even underperforming RISE. We also observed that during training, 3DTacDex-P achieves lower hand-joint prediction error compared to other methods. These are likely due to overfitting to tactile patterns, which are similar across different expert demonstrations, while visuals varied a lot due to the larger workspace compared to the original method. The FoAR baseline, despite careful threshold selection for contact prediction, is unstable. While it performs well on the Reorientation task, it fails on both Open Box and Flip tasks. In these failures, the policy reaches the object but cannot apply the correct manipulation forces. Expert data reveals frequent contact changes, making manual thresholding unreliable and leading to inconsistent contact labels. In contrast, Ours mitigates tactile overfitting by dynamically adjusting visuo-tactile attention across diverse tasks without relying on manual labeling, demonstrating robust generalization and consistent performance.
@@ V-D Effectiveness of Force-Guided Attention Fusion
V-E Importance of Future Force Prediction and Guidance
As shown in Tab. III, without both FFP and FFG, the success rate is only 50%. Incorporating observed force prediction (OFP) and guidance (OFG) increases the success rate to 70%, demonstrating the effectiveness of force prediction and guidance. Replacing observed force (OFP) with future force prediction (FFP) further raises the success rate to 90%, similar to Ours, highlighting the importance of predicting future force. We also observe that for both OFP and FFP, more attention is given to the tactile modality upon contact, indicating that force prediction effectively reinforces tactile modality and improves balance. However, Ours w/o FFG often requires multiple attempts, leading to an AEL of 166, whereas Ours typically succeeds on the first try. Additionally, Ours w/o FFG exhibits more risky behaviors, such as continuously squeezing the board. These findings underscore the importance of FFG.
@@ V-F Analysis of Attention
We visualize the attention weights across different steps, as shown in Fig. 3. For all tasks, during the reaching stages, more attention is given to the visual, as the hand manipulates the object, the attention shifts towards the tactile, indicating effective attention adjustment. However, we observe that the changes in attention weights for the reorientation task are not substantial. Tactile features indeed gain more attention during contact, but visual attention remains relatively high. This is likely due to the importance of visual features not only during reaching but also throughout manipulation, as tasks like reorientation require vision to assess whether the object has been rotated to the correct angle.
@@ V-G Generalization on Unseen Objects
We validate the generalization of our method by testing it on five objects with varying colors and geometries, as shown in Fig. 4. Each object is tested four times with different random poses. As shown in Tab. IV, our policy achieves a 75% success rate on unseen objects, demonstrating strong generalization, even for objects with significantly different geometries, such as a small paper cup in the reorientation task and a whiteboard eraser in the flipping task.
@@
Ours 85% 75% 65% 75%
@@ VI CONCLUSIONS
VII Acknowledgments
We thank Hongjie Fang for insightful discussions, Tianshu Wu for assistance with camera calibration, and Mingjie Pan and Gu Zhang for discussions on point cloud networks. We also appreciate valuable feedback on the draft from Yujie Zhao, Hongwei Fan, Zhiyuan Ma, and Qiyang Yan, as well as the RISE/FoAR authors for their released code. This project was supported by the National Youth Talent Support Program (8200800081), National Natural Science Foundation of China (62376006) and National Natural Science Foundation of China (62136001).